### v2

# Creation of domain-wall Skyrmions in chiral magnets

Abstract
We study the capture of a magnetic Skyrmion into a domain wall (DW) structure in chiral magnets and find that in the respective ground states of the DW and the Skyrmion, they repel each other. This means that an isolated magnetic Skyrmion cannot a priori enter the DW and become a DW-Skyrmion. However, rotating the DW’s phase away from the stable phase may cause the successful capture and hence creation of a stable bound state of a magnetic Skyrmion and a DW: the DW-Skyrmion. At certain distances, the DW may also destroy the magnetic Skyrmion by inducing shrinkage of the latter. This happens as the isolated magnetic Skyrmion has negative DMI energy, whereas the DW-Skyrmion has positive DMI energy; in a finite range from the DW, the magnetic Skyrmion can be pushed to have vanishing DMI energy, for which it collapses to a point. En passant, we find the possibility of pair creation of a Skyrmion anti-DW-Skyrmion pair and a creation of more than one DW-Skyrmions by a Kibble-like mechanism.
1 Introduction
Skyrmions are particle-like topological solitons supported by the third homotopy group $\pi_{3}(S^{3})\simeq{\mathbb{Z}}$ in the Skyrme model [1, 2, 3] in which baryons are described as Skyrmions. The Skyrme model is justified in the large-$N_{c}$ limit (large number of colors) to characterize nucleons properly [4, 5, 6]. When Skyrmions coexist with a domain wall (DW), the former are expected to be absorbed into the latter, forming a composite state called a DW-Skyrmion [7, 8, 9, 10, 11, 12]. Recently, they have been found to support a new phase, the DW-Skyrmion phase in QCD, in the presence of strong magnetic field or rapid rotation [13, 14, 15, 16, 17, 18].
Baby Skyrmions are two-dimensional analogues of Skyrmions supported by the second homotopy group $\pi_{2}(S^{2})\simeq{\mathbb{Z}}$ instead of the third one [19, 20]. Recently, magnetic Skyrmions have attracted great attention in condensed matter physics. Magnetic Skyrmions are particle-like solutions in the magnetization vector of two-dimensional chiral magnetic materials [21, 22, 23]. The magnetic Skyrmion owes its stability to the Dzyaloshinskii-Moriya interaction (DMI) that is induced by the spin-orbit coupling [24, 25], in contrast to the baby Skyrmion which is stabilized by the four derivative term, known as the Skyrme term. Magnetic Skyrmions are important candidates for future storage systems in the form of race track memory, that potentially consumes much less energy than current data storage technology [26], for a review see [27]. Magnetic Skyrmions were first realized experimentally in chiral magnetic materials [28, 29]. Chiral magnets contain also other extended solutions that are string-like, called magnetic DWs (domain lines) that separate magnetic domains. In particular, in the plane there exists domain lines [30, 31]. Both the DW (or domain line in 2D) and the magnetic Skyrmion are topologically protected objects, but the DW is stable and exists simply in the presence of the easy-axis anisotropy. Both the Skyrmion and the DW can be of two types: Bloch or Néel, depending on how the magnetization vector twists around when going from the center of the Skyrmion to its outside, whereas the magnetization vector on a Bloch (Néel) DW is parallel (perpendicular) to the DW. In 3-dimensional samples the magnetic domain wall is a surface object, and the magnetic Skyrmion extends as a (twisted) string-like object through the material which can even form braids [32]. One of the difficulties in controlling Skyrmions under an applied current is the Skyrmion Hall effect [33, 34, 35]: the Skyrmion trajectory bends. To overcome this problem, it may be useful to consider Skyrmions trapped inside a DW where Skyrmions are restricted to move along the domain line. DW-Skyrmions in two dimensions were proposed in a field theoretical model [36, 37, 38, 39]. In the planar samples of chiral magnets, domain lines can stably host a Skyrmion [40, 41, 42, 43, 44, 45, 46, 47, 48, 49], which creates a little cusp on the line. This object has been coined a magnetic DW-Skyrmion and has been discovered in the laboratory [50, 51, 52]. The domain-wall Skyrmions in chiral soliton lattices are unstable to decay into two merons [48], which are quasi-particles with topological charge $1/2$ [53, 54, 55, 56].
In this paper, we find that capturing an isolated (stable) magnetic Skyrmion into a (stable) DW is not as easy as anticipated. Indeed, we confirm with an exact analytic calculation of the asymptotic interactions as well as with detailed numerical computations, that the isolated Skyrmion in the bulk repels the DW. In order to capture the isolated Skyrmion into the DW, we propose two ways of altering the solitons so as to make a successful capture possible: We change the complex phase of the DW or rotate the Skyrmion in the target space (2-dimensional isorotation). By rotating the DW’s phase into the unstable regime, we find that a successful capture can take place on separation distances of up to about 4.5 times the Skyrmion radius. Rotating the Skyrmion, on the other hand, requires the Skyrmion to be in very close proximity to the DW, in order for the capture to be successful. We also find that the stable Skyrmion becomes unstable when placed too close to a stable DW. Finally, we discover a Kibble-like mechanism that can create both anti-DW-Skyrmions as well as multiple-DW-Skyrmion at the complex phase where the DW is maximally unstable.
This paper is organized as follows. In sec. 2, we introduce the magnetic Skyrmion model with the anisotropy potential. In sec. 2.1, we introduce the setup which is used as the initial condition for numerical computations and present the possible final states. In sec. 2.2, we present the numerical computations that lead to the phase diagram from which one can read off where a successful capture of an isolated Skyrmion into a DW is possible. In sec. 2.3, we compute the asymptotic interactions between the isolated Skyrmion and the DW. Finally, we conclude in Sec. 3 with a discussion and outlook. We have relegated the connection between Bloch and Néel type DWs and Skyrmions to appendix A.
2 The chiral magnetic model
The energy functional for a chiral magnetic material is given by
$E=\int\mathcal{E}\mathop{}\!\mathrm{d}^{2}x,\qquad\mathcal{E}=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\kappa\mathbf{n}\cdot\nabla\times\mathbf{n}+V(\mathbf{n}),$ (2.1)
with the anisotropy term
$V(\mathbf{n})=\frac{1}{2}(1-n_{3}^{2}),$ (2.2)
where we have rescaled length and energy units so as to only have dependence on one physical parameter, $\kappa$ [^1] The physical constants of ref. [57] correspond to the value of the coupling $\kappa\approx 0.4$. The field $n$ is interpreted as the magnetization vector in chiral magnets, but here we take it as a nonlinear sigma model field with target space $S^{2}$: $\mathbf{n}=(n_{1},n_{2},n_{3})$ satisfying $\mathbf{n}\cdot\mathbf{n}=1$.
The magnetic Skyrmion carries a topological charge given by
$Q=\int\mathcal{Q}\;\mathop{}\!\mathrm{d}^{2}x,\qquad\mathcal{Q}=-\frac{1}{4\pi}\mathbf{n}\cdot\partial_{1}\mathbf{n}\times\partial_{2}\mathbf{n},$ (2.3)
which arises by considering the stereographic projection of the plane to the 2-sphere, so that the magnetization vector is a map from the 2-sphere (the plane) to a unit vector in $\mathbb{R}^{3}$, i.e. another 2-sphere: this is characterized by $\pi_{2}(S^{2})=\mathbb{Z}\ni Q$, with $Q$ being the topological charge given above.
The equation of motion corresponding to the energy (2.1) is
$\partial_{i}^{2}n^{a}-2\kappa\epsilon^{aib}\partial_{i}n^{b}+n^{3}\delta^{a3}-\Lambda n^{a}=0,$ (2.4)
where $\Lambda$ is a Lagrange multiplier used to impose the nonlinear sigma model constraint $\mathbf{n}\cdot\mathbf{n}=1$. Changing to stereographic coordinates, $u$, the equation of motion becomes
$\mathrm{eom}$ $:=(1+|u|^{2})\partial_{i}^{2}u-2\bar{u}(\partial_{i}u)^{2}-2\kappa\left[\mathrm{i}(u+\bar{u})\partial_{1}u+(u-\bar{u})\partial_{2}u\right]-u(1-|u|^{2})=0,$ (2.5)
where $u$ is defined as
$u=\frac{n_{1}+\mathrm{i}n_{2}}{1+n_{3}},$ (2.6)
which is defined everywhere on $S^{2}$, except at the south pole ($n_{3}=-1$).
2.1 Setup and initial state
First, we will construct the composite initial configuration made of a magnetic Skyrmion placed at the origin and a domain wall sitting at $x=X_{0}<0$ with its world-line pointing in the $\hat{y}$-direction, see fig. 1.
$y$$x$$|X_{0}|$$(0,0)$
Fig. 1: Setup of DW and isolated Skyrmion as initial condition (2.7).
Each of the two solitons are solutions to the equations of motion if they are well separated and the initial configuration is made using the superposition rule in stereographic coordinates
$u^{\rm composite}=u^{\rm sk}+u^{\rm DW},$ (2.7)
where the stereographic coordinate is defined in eq. (2.6). The magnetic Skyrmion and domain wall are given by
$u^{\rm sk}=e^{\mathrm{i}(\theta+\beta)}\tan\left(\frac{f(r)}{2}\right),\qquad u^{\rm DW}=e^{\mathrm{i}\alpha-(x-X_{0})},$ (2.8)
where the coordinate system in the plane is $x+\mathrm{i}y=re^{\mathrm{i}\theta}$ with the Skyrmion placed at the origin and the domain wall at $x=X_{0}$ parallel to the $\hat{y}$-direction. We take $\kappa>0$ for which the stable magnetic Skyrmion and the stable DW correspond to $\beta=\tfrac{\pi}{2}$ and $\alpha=\tfrac{\pi}{2}$, respectively. The sign of the DW is chosen such that to the left of the DW ($x<X_{0}$) the magnetization in the ground state is $n_{3}=-1$ while to the right of the DW ($x>X_{0}$) $n_{3}=1$. The well-separated solitons have energies
$E^{\rm sk}$ $=\int\bigg{[}(f^{\prime})^{2}+\sin^{2}f\left(1+\frac{1}{r^{2}}+4\kappa\sin\beta f^{\prime}\right)\bigg{]}\pi r\mathop{}\!\mathrm{d}r,$
$E^{\rm DW}$ $=\int\left[2-\pi\kappa\sin\alpha\right]\mathop{}\!\mathrm{d}y,$ (2.9)
from which it is clear that $\alpha=\frac{\pi}{2}$ is the minimal energy for the DW, while for the Skyrmion $\beta=\frac{\pi}{2}$ is the minimal energy too, since $f^{\prime}=\frac{\mathop{}\!\mathrm{d}f}{\mathop{}\!\mathrm{d}r}$ is negative. The DW solution ($u^{\rm DW}$) in eq. (2.8) is exactly the renowned sine-Gordon solution in the inhomogeneous $\mathbb{C}P^{1}$ coordinate.
Inserting the Skyrmion field, $u^{\rm sk}$, of eq. (2.8) into the equation of motion (2.5) yields
$f^{\prime\prime}+\frac{f^{\prime}}{r}-\frac{\sin 2f}{2r^{2}}+\frac{2\kappa\sin\beta\sin^{2}f}{r}-\frac{\sin 2f}{2}=0,$ (2.10)
$\kappa\cos\beta\sin(f)f^{\prime}=0.$ (2.11)
The first and the second equation correspond to the real and imaginary part of the complex equation (2.5), respectively. The imaginary part fixes $\beta=\pm\frac{\pi}{2}$, with the upper sign being the stable phase of the Skyrmion and the lower sign the unstable phase. Choosing the stable phase, $\sin\beta=1$ in eq. (2.10). Inserting instead the DW field, $u^{\rm DW}$, of eq. (2.8) into the equation of motion (2.5) yields
$\kappa e^{-\mathrm{i}\alpha-2(x-X_{0})}\cos\alpha=0,$ (2.12)
which fixes the DW phase to be $\alpha=\pm\frac{\pi}{2}$, with the upper sign being the stable phase of the DW and the lower sign the unstable phase, as can also be seen from eq. (2.9).
Note that the isolated Skyrmion is unstable for $\kappa\sin\beta<0$, which renders the DMI energy contribution positive. This is a consequence of Derrick’s theorem [58]; Derrick’s theorem states that solitons collapse in more than one spatial dimensions in a theory with a kinetic (gradient) term and a potential. The DMI term is a loophole, since it is not classically conformal invariant (like the kinetic term) and it is negative:
$E^{\lambda}=E_{2}+\lambda E_{\rm DMI}+\lambda^{2}E_{0},$ (2.13)
where $E_{2}$ is the kinetic (gradient) term, $E_{\rm DMI}$ is the DMI term, $E_{0}=\int V\mathop{}\!\mathrm{d}^{2}x$ is the potential term and $x\to\lambda x$ is a scale transformation. For negative DMI energy, stabilization of the Skyrmion obeys the virial law ($\partial E^{\lambda}/\partial\lambda|_{\lambda=1}=0$):
$E_{\rm DMI}=-2E_{0},$ (2.14)
where $E_{0}>0$. This is only possible for when the DMI energy is negative and hence when $\beta=\pi/2$ in our notation.
We should emphasize that the superposition or “Abrikosov Ansatz” used in eq. (2.7) is not a solution, but simply an initial guess from which we will start numerical computations. At asymptotic distances, however, the superposition becomes a very good approximation. It is worth to point out that the superposition of two solitons in the inhomogeneous coordinates $u$ is very advantageous, since the composite soliton already satisfies the nonlinear constraint of the nonlinear sigma model via eq. (2.6).
2.2 Numerical results
We will now turn to numerical computations. In order to explore the possibility of a successful capture of an isolated Skyrmion into the DW, we perform a large-scale set of simulations using the initial condition (2.7) and evolving it to the nearest final state, using the so-called arrested Newton flow algorithm. The simulations are performed using a finite-distance approximation to derivatives using a fourth-order 5-point stencil on a square lattice of $682^{2}$ lattice sites. The arrested Newton flow method is a second-order fictitious time-evolution (with fictitious time $\tau$) of the equations of motion (i.e. akin to a wave equation)
$\partial_{\tau}^{2}u=\textrm{eom},$ (2.15)
with the special condition that the kinetic energy is removed (i.e. $\partial_{\tau}u=0$) every time the potential part of the energy (i.e. (2.1)) increases and eom is given in eq. (2.5). The Newton flow accelerates towards the nearest local minimum of the energy functional (2.1) and if not for the conditional arrest, it would blast right past it. Arresting it every time the flow has just passed the minimum ensures a fast convergence to the nearest local minimum. For simplicity, we use Neumann boundary conditions on the edges of the lattice, which allows for the topological charges to enter or exit the simulation area.
2.2.1 The unsuccessful capture
Naturally, as the first attempt we place the stable isolated Skyrmion (i.e. with $\beta=\pi/2$) near the stable DW (i.e. with $\alpha=\pi/2$); two outcomes can occur, either the Skyrmion is repelled (for large enough separation distance) or it is annihilated (for short enough separation distance). This is bad news for the capture and creation of a DW-Skyrmion starting from an isolated Skyrmion and an empty DW. We will get back to why the Skyrmion is annihilated by the DW shortly.
2.2.2 Phase diagram and final states
In order to search for a successful capture, we perform a scan over the phase parameters $\alpha$ and $\beta$ for a range of separation distances $X_{0}$. The phase diagram computed by running one simulation for each point in a lattice of initial condition parameters with 64 points parametrizing the angle $\alpha$ ($\beta$) and $74$ points parametrizing the separation distance in the range $X_{0}\in[-6.2,-R]$, with $R\approx 0.92$ being the Skyrmion radius. This yields a grand total of $4736$ simulation for the $\alpha$-phase diagram with $\beta=\frac{\pi}{2}$ and the same number for varying $\beta$ at fixed $\alpha=\frac{\pi}{2}$. Due to the large number of simulations, each with a large number of flops, we used our CUDA C code on a GPU cluster. The final states are shown in fig. 2 and the resulting phase diagram, displaying which final state is reached by the numerical computation for the given values of $\alpha$, $\beta=\frac{\pi}{2}$ and $X_{0}$ ($\alpha=\frac{\pi}{2}$, $\beta$ and $X_{0}$) are shown in fig. 3a (fig. 3b).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption $\qquad\mathcal{E}\qquad\qquad\qquad\qquad$ $\mathcal{E}_{\rm DMI}\qquad\qquad\qquad$ $\mathcal{Q}\qquad\qquad\qquad\qquad$ $\mathbf{n}$
Fig. 2: The energy density ($\mathcal{E}$), the DMI energy contribution to the energy ($\mathcal{E}_{\rm DMI}$), the topological charge density ($\mathcal{Q}$) and the magnetization vector $\mathbf{n}$, for the five color-coded final states (colors shown left of each row): DW + Skyrmion (blue), DW-Skyrmion (green), only DW (red), anti-DW-Skyrmion and Skyrmion pair that annihilates and leaves behind two DW-Skyrmions (yellow) and finally the same as the previous case with the anti-DW-Skyrmion and bulk Skyrmion being a metastable bound state (magenta). These colors are referenced in fig. 3. The colors in the magnetization vector plot, $\mathbf{n}$, (fourth column) are defined as: $n_{3}=1$ is white, $n_{3}=-1$ is black, $n_{2}=1$ is green, $n_{1}=1$ is red, $n_{2}=-1$ is magenta and $n_{1}=-1$ is cyan.
Refer to caption Refer to caption
Fig. 3: The phase diagram of final states as a function of the initial parameters (a) $(X_{0},\alpha)$ for $\beta=\frac{\pi}{2}$ and (b) $(X_{0},\beta)$ for $\alpha=\frac{\pi}{2}$. The states are color coded with the same colors as in fig. 2: DW + Skyrmion (blue), DW-Skyrmion (green), only DW (red), anti-DW-Skyrmion and Skyrmion pair that annihilates and leaves behind two DW-Skyrmions (yellow) and finally the same as the previous case with the anti-DW-Skyrmion and bulk Skyrmion being a metastable bound state (magenta). The magenta state eventually decays into the yellow state and the yellow state can decay into the red state for finite specimen sizes. The marked points with (e.g. A1,B1$\cdots$) correspond to videos available in the supplemental material.
The initial condition $u^{\rm composite}$ is in general unstable and each numerical computation will flow the specific initial configuration with $(\alpha,\beta,X_{0})$ to its nearest local minimum of the energy functional, which we shall denote final states. This results in 5 different final states, which we show in fig. 2 (of which 2 are metastable):
  • the separated magnetic Skyrmion and DW with increasing separation distance (they repel each other in their respective stable stable states (blue color code),
  • the DW-Skyrmion where the magnetic Skyrmion has been captured into the world volume of the DW (green color code),
  • the annihilation of the magnetic Skyrmion by shrinkage (red color code),
  • a metastable state with an anti-Skyrmion induced on the DW by the long-distance forces from a Skyrmion in the bulk, which quickly annihilates and leaves a pair of DW-Skyrmions on the DW (yellow color code),
  • same as the above-described state, but with the anti-DW-Skyrmion and the bulk Skyrmion being a long-lived composite before the eventual annihilation (magenta color code).
Starting with the stable Skyrmion ($\beta=\frac{\pi}{2}$), we obtain, as promised, that the stable DW ($\alpha=\frac{\pi}{2}$) annihilates the Skyrmion for separation distances smaller than $|X_{0}|\lesssim 1.75$ whereas the Skyrmion is repelled by the DW for larger separation distances. In order to capture the isolated stable Skyrmion into the DW for the creation of the DW-Skyrmion composite soliton, one approach that works and is found for the first time in this paper, is by modifying the phase of the DW so that it is not in its stable phase (ground state). This corresponds to either increasing or decreasing $\alpha$ from $\alpha=\frac{\pi}{2}$. The unstable DW phase is not enough by itself, as the DW will simply flow back to its stable position (ground state); hence, it is also necessary that the Skyrmion is in close proximity to the DW at the time when the DW phase is altered from its stable value. This capture corresponds to the green-colored area in fig. 3a, which forms a wedge in both directions away from the stable $\alpha=\frac{\pi}{2}$ phase. The longest distance for which the Skyrmion can be captured by the DW is found near the most unstable phase of the DW, i.e. $\alpha=\frac{3\pi}{2}$ (with the exception for the exact antipodal value $\alpha=\frac{3\pi}{2}$ where an array of other phenomena appears, see below).
Turning now to the most unstable point of the DW, i.e. $\alpha=\frac{3\pi}{2}$, the DW is now on the top of the hill in the static energy landscape and has the potential to create Skyrmion-anti-Skyrmion pairs. This is simply the 1-dimensional version of the Kibble mechanism [59, 60]. Now, instead of random perturbations, as in the cosmological setting of the Kibble mechanism, here the perturbation is specific and is given by the interaction between the DW and the (isolated) Skyrmion. For separation distances shorter than $|X_{0}|\lesssim 5$, the perturbation of the unstable DW creates an anti-Skyrmion on the DW, but the anti-DW-Skyrmion is in such close proximity to the isolated Skyrmion in the bulk, that the two solitons annihilate and disappear. In the process of perturbing the DW, the perturbation also creates a pair of DW-Skyrmions on each side of the anti-DW-Skyrmion. The minimization of energy, however, pushes them away and off of our finite-sized computation grid. Increasing the separation distance ($|X_{0}|$), at the unstable DW phase (i.e. $\alpha=\frac{3\pi}{2}$), still creates an anti-DW-Skyrmion by perturbing the unstable DW, but now the bulk (isolated) Skyrmion is so far away that they remain as a metastable bound state for some time – until they eventually come close enough to each other to annihilate.
We now turn to the more physically difficult case of leaving the DW in its stable phase and rotating the isolated Skyrmion (in target space coordinates). In this case, the isolated Skyrmion becomes unstable when the phase is $\beta=\frac{3\pi}{2}$ for which the DMI energy is positive and hence there is no loophole to avoiding Derrick’s theorem. When the (isolated) Skyrmion is sufficiently far from the DW, indeed this is what we see: the Skyrmion collapses and disappears (see the red stripe at large $|X_{0}|$ at $\beta=\frac{3\pi}{2}$ in fig. 3b). We already know that for a stable (isolated) Skyrmion the (stable) DW repels or annihilates the Skyrmion, so the only possibility of capture is when the Skyrmion is rotated by some angle away from $\beta=\frac{\pi}{2}$. Indeed, this is what we see from fig. 3b. We find, however, that only when the (isolated) Skyrmion is in very close proximity to the DW, a successful capture is possible, see the green-shaded area in fig. 3b.
It is clear that when the isolated Skyrmion in the bulk is rotated into the unstable phase ($\beta=\frac{3\pi}{2}$), it decays in the absence of interaction with the DW, see the large-$|X_{0}|$ (left) part of fig. 3b.
$\ \ \mathbf{n}$Refer to caption Refer to caption Refer to caption $\mathcal{E}_{\rm DMI}$efer to caption Refer to caption Refer to caption
Fig. 4: Shrinkage of a Skyrmion in the near proximity to the DW. The left column corresponds to the initial configuration (2.7) (with $X_{0}=-1.5$), the middle to arrested Newton flow (aNF) step number 150 and the right-most column to step number 250. In this figure $\alpha=\beta=\frac{\pi}{2}$ and $\kappa=0.4$.
The collapse of the stable (isolated) Skyrmion (i.e. with $\beta=\frac{\pi}{2}$) at close proximity to the DW can be explained as follows. The DMI energy of the isolated Skyrmion is negative (see the second column of the first row (blue) in fig. 2), whereas the DMI energy of the DW-Skyrmion is positive (see the second column of the second row (green) in fig. 2). When there is enough “momentum” to overcome the barrier from going from an isolated bulk Skyrmion to becoming a DW-Skyrmion, it works out. If there the Skyrmion gets “stuck” midway, it suffers the issue of being repelled and gaining a negative DMI energy or being attracted and gaining a positive DMI energy. With a nearly vanishing DMI energy, there is no loophole to Derrick’s theorem and the soliton collapses. Hence the demise of the Skyrmion, see fig. 4.
The final issue that needs more theoretical explanation, is why does a stable (isolated) Skyrmion and a stable DW repel each other? We will explain this using an asymptotic analytic calculation in the next subsection.
2.3 Interactions
The aim of this section is to understand, qualitatively, why the stable DW [$\alpha=\frac{\pi}{2}$] and the stable (isolated) Skyrmion [$\beta=\frac{\pi}{2}$] repel each other at large separation distances. This can be understood by studying the asymptotic interactions.
The analytic computation of the interaction energy is found as follows. We separate the plane into two regions: Left and right and place the DW on the left and the magnetic Skyrmion on the right. Each field is assumed to be small in the other soliton’s core region and we assume the soliton in its own core region to obey the equation of motion. The interaction energy is defined as the total energy with each of the individual solitons’ energies subtracted off
$E^{\rm int}=E^{\rm DW+sk}-E^{\rm DW}-E^{\rm sk}.$ (2.16)
Expanding the energy (2.1) around a solution as $u\to u+\delta u$, the expression for the interaction energy will simplify as the part of the energy of the soliton (DW or Skyrmion) will cancel out. In order to perform the computation, however, it is important that the variational problem is well defined, which as described in refs. [61, 62, 63] requires that we drop a boundary term in the energy functional. The equation of motion is of course unchanged by dropping this boundary term. Integrating the energy (2.1) by parts, we identify the boundary term that should be dropped
$\mathcal{E}$ $=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\kappa\epsilon^{aib}n^{a}\partial_{i}n^{b}+\frac{1}{2}[1-(n^{3})^{2}]$
$=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\kappa\sum_{a=1}^{2}\epsilon^{aib}n^{a}\partial_{i}n^{b}-\kappa\epsilon^{ib3}\partial_{i}n^{3}n^{b}+\frac{1}{2}\left[1-(n^{3})^{2}\right]+\kappa\epsilon^{ib3}\partial_{i}(n^{3}n^{b}).$ (2.17)
Expanding the energy functional without the boundary term (i.e. the last term) about the perturbations, we obtain
$\mathcal{E}$ $=\mathcal{E}_{0}-\frac{2\,\overline{\mathrm{eom}}}{(1+|u|^{2})^{3}}\delta u+\partial_{i}\left(\left[\frac{2\partial_{i}\bar{u}}{(1+|u|^{2})^{2}}-\frac{4\mathrm{i}\kappa(\bar{v}_{i}u-v_{i}\bar{u})}{(1+|u|^{2})^{3}}\right]\delta u\right)+\textrm{c.c.}+\mathcal{O}(\delta u^{2},\delta\bar{u}^{2}).$ (2.18)
The equation of motion (eom) vanishes on the background solution. Hence, to linear order we obtain [^2] Changing the phase $\alpha$ for the DW or the phase $\beta$ for the Skyrmion away from the stable point $(\alpha=\beta=\frac{\pi}{2})$ corresponds to turning on a quadratic field operator in the piece eom, which we neglect to leading order.
$\mathcal{E}=\mathcal{E}_{0}+2\partial_{i}(\partial_{i}\bar{u}\delta u)+2\partial_{i}(\partial_{i}u\delta\bar{u})+\mathcal{O}(\delta u^{2},\delta\bar{u}^{2}).$ (2.19)
Specializing now to the case of the DW sitting on the left and the Skyrmion on the right, we may write the interaction energy as
$E^{\rm int}$ $=2\int_{L}\partial_{i}\left[\partial_{i}u^{L}\bar{u}^{R}+\textrm{c.c.}\right]\mathop{}\!\mathrm{d}^{2}x+2\int_{R}\partial_{i}\left[\partial_{i}u^{R}\bar{u}^{L}+\textrm{c.c.}\right]\mathop{}\!\mathrm{d}^{2}x$
$=2\int_{\Gamma}\left[\partial_{i}u^{L}\bar{u}^{R}-\partial_{i}u^{R}\bar{u}^{L}+\textrm{c.c.}\right]\mathop{}\!\mathrm{d}S^{i},$ (2.20)
where $\Gamma$ is a line in the $\hat{y}$-direction, separating the DW and the Skyrmion. In the left region, we use $u=u^{L}=u^{\rm DW}$ as the background solution with $\delta u=u^{R}=u^{\rm sk}$ being the perturbation and in the right region, we use $u=u^{R}=u^{\rm sk}$ as the background solution with $\delta u=u^{L}=u^{\rm DW}$ being the perturbation. We assume that both $u^{L}=u^{\rm DW}$ and $u^{R}=u^{\rm sk}$ are small on the line $\Gamma$ that separates the left and the right regions.
In order to evaluate the line integral in eq. (2.20), we use Green’s theorem and turn the integral into an area integral over the right-hand domain, that includes the Skyrmion
$E^{\rm int}$ $=-2\int_{R}\left[\Delta u^{L}\bar{u}^{R}-\Delta u^{R}\bar{u}^{L}+\textrm{c.c.}\right]\mathop{}\!\mathrm{d}^{2}x.$ (2.21)
Since this is a mathematical trick to evaluate the line integral on $\Gamma$, we can now use that both $u^{L}$ and $u^{R}$ can be described by their exact solutions to their respective linearized equations of motion. Since $u^{L}=u^{\rm DW}$ is already exact and exponentially small in the right region, this applies to the Skyrmion field $u^{R}=u^{\rm sk}$. Linearizing the equation of motion for the Skyrmion (2.10), we get
$f^{\prime\prime}+\frac{f^{\prime}}{r}-\frac{f}{r^{2}}-f=0,$ (2.22)
which has the exact solution $f=qK_{1}(r)$, and in turn
$u^{\rm sk,\,linear}\simeq\frac{q}{2}e^{\mathrm{i}(\theta+\beta)}K_{1}(r)=-qe^{\mathrm{i}\beta}\partial_{\bar{z}}K_{0}(r),$ (2.23)
where $K_{1}(r)$ and $K_{0}(r)$ are modified Bessel functions of the second kind. Comparing with the 2D Green function $(\Delta-1)K_{0}(r)=-2\pi\delta^{(2)}(\mathbf{x})$, we get that the equation for the asymptotic solution $u^{\rm sk}$ is
$(\Delta-1)u^{R}=2\pi qe^{\mathrm{i}\beta}\partial_{\bar{z}}\delta^{(2)}(\mathbf{x}),$ (2.24)
which has a delta-function source at its origin [19, 64]. This linearized solution is valid on the line $\Gamma$ and the line integral is by Green’s theorem equal to the area integral using this delta-function sourced linearized solution. The DW field, on the other hand, is not sourced in the right region, hence
$(\Delta-1)u^{L}=0.$ (2.25)
Using eqs.  (2.24) and (2.25), we obtain
$E^{\rm int}$ $=2\int_{R}\left[\bar{u}^{L}(\Delta-1)u^{R}+\textrm{c.c.}\right]\mathop{}\!\mathrm{d}^{2}x$
$=4\pi q\cos(\beta-\alpha)e^{-|X_{0}|},$ (2.26)
with $q$ being a positive constant, $\alpha$ and $\beta$ are the phases of the DW and the Skyrmion, respectively, and finally $X_{0}(<0)$ is the DW position. We see from the analytic leading-order result, that the stable Skyrmion and stable DW ($\beta=\alpha=\frac{\pi}{2}$) yield a positive interaction energy, corresponding to repulsion. On the other hand, if either the DW or the Skyrmion is rotated away from their respective stable phases, attraction is possible. This explains why the well separated stable DW and stable (isolated) Skyrmion repel each other. At short distances, this asymptotic interaction cannot be trusted as nonlinear terms become important for the minute details of the interactions.
3 Conclusion and outlook
In this paper, we have investigated the capture of a single isolated Skyrmion into a stable DW in a chiral magnet. It turns out that they repel each other at large distances and at short distance the Skyrmion is obliterated by the following effect: The stable isolated Skyrmion in the bulk has a negative DMI energy, whereas the stable DW-Skyrmion has a positive DMI energy. At short distances, if the Skyrmion does not have a momentum to overcome a certain (nonlinear) barrier, the DMI energy remains small or positive leading to a shrinking instability of the Skyrmion, unless the Skyrmion manages to be pushed all the way into the DW and become a stable DW-Skyrmion – a cuspy kink of twisted spins as a sine-Gordon soliton living on the DW. In this paper we propose two ways to obtain a successful capture of the isolated Skyrmion in the bulk into the DW: Either we perturb the DW into an unstable complex phase (away from its minimum of the energy) or we rotate the Skyrmion on the target space, which increases its DMI energy (from a negative value towards zero). Both perturbations can lead to a successful capture of the isolated Skyrmion into the DW, provided that the Skyrmion is in close enough proximity to the DW. This is mapped out in the phase diagram shown in fig. 3. When the DW is perturbed to the maximally unstable complex phase ($\alpha=\frac{3\pi}{2}$), it can fall into the stable phase ($\alpha=\frac{\pi}{2}$) in two different directions. This is akin to the 1-dimensional version of the Kibble mechanism, which indeed can create DW-Skyrmions or anti-DW-Skyrmions. As opposed to the Kibble mechanism, the perturbations are not random, but are indeed given by the asymptotic interaction between the DW and the Skyrmion, which creates an anti-DW-Skyrmion on the DW facing the bulk Skyrmion. This pair of solitons annihilate and disappear, except for a large enough separation distance, where they appear as a metastable bound state for quite some time, before they eventually collide and annihilate. Finally, on the theoretical side we show analytically that the asymptotic interaction between an isolated Skyrmion and a DW leads to a repulsive force.
With our proposals, we have found two ways to capture an isolated Skyrmion into a DW, i.e. by rotating the DW phase or by rotating the Skyrmion on the target space, but there could be many more ways to perform a successful capture – some of all these could become experimentally viable in real chiral magnets. Although we performed the analysis using the Bloch DMI by convention, we could equivalently have used a Néel DM instead (see Appendix A).
A future direction of study would be to use the Landau-Lifshitz-Gilbert (LLG) equation instead of energy minimization (here arrested Newton flow). An LLG flow without a current turned on corresponds to a gradient flow, which is essentially an exponentially slower version of the arrested Newton flow, but with an added symplectic flow that pushes the “particles” in the flow in the orthogonal direction to the flow. For a field and hence a soliton, this has consequences modifying the phase diagram of the final states. We expect shifts in the phase diagram in the angular directions (i.e. $\alpha$ or $\beta$), but otherwise qualitatively similar behaviors up to angular shifts. Another interesting direction for future studies would be to find a way to start with a DW-Skyrmion and somehow eject the Skyrmion so that one ends up with an empty DW and an isolated Skyrmion – viz. the inverse process of what we have studied in this paper.
Note added
After posting this paper on the arXiv, a preprint with some overlap also appeared on the arXiv [65].
Acknowledgments
S. B. G. thanks the Outstanding Talent Program of Henan University for partial support. This work is supported in part by JSPS KAKENHI [Grants No. JP23KJ1881 (Y. A.), No. JP22H01221 and No. JP23K22492 (M. N.)] and the WPI program “Sustainability with Knotted Chiral Meta Matter (WPI-SKCM2)” at Hiroshima University (M. N.).
Appendix A Bloch versus Néel DW-Skyrmion
We can formally consider the static energy of the form of a gauged nonlinear sigma model [66, 67, 46]
$\mathcal{E}$ $=\frac{1}{2}D_{i}\mathbf{n}\cdot D_{i}\mathbf{n}+m^{2}\big{(}1-(n^{3})^{2}\big{)}$
$=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\frac{1}{2}\mathbf{A}_{i}\cdot\mathbf{A}_{i}-\frac{1}{2}(\mathbf{A}_{i}\cdot\mathbf{n})^{2}+\partial_{i}\mathbf{n}\cdot\mathbf{A}_{i}\times\mathbf{n}+m^{2}\big{(}1-(n^{3})^{2}\big{)},$ (A.1)
with the gauge covariant derivative
$D_{i}\mathbf{n}=\partial_{i}\mathbf{n}+\mathbf{A}_{i}\times\mathbf{n},$ (A.2)
and $\mathbf{n}=(n^{1},n^{2},n^{3})$ is the magnetization field as usual. We choose a constant background gauge field ($\mathbf{A}_{i}=(A_{i}^{1},A_{i}^{2},A_{i}^{3})$) as
$A_{i}^{a}=\begin{cases}-\kappa\delta_{i}^{a},\qquad&\textrm{Dresselhaus SOC},\\ -\kappa\epsilon_{i}^{\phantom{i}a},&\textrm{Rashba SOC}.\end{cases}$ (A.3)
In both cases, the field strength is
$F_{12}=F_{12}^{a}\frac{\tau^{a}}{2}=\kappa^{2}\tau^{3},$ (A.4)
implying that the above two choices are transformed to each other by a gauge transformation, given below. Inserting the two choices of gauge field into the energy functional (A.1) yields
$\mathcal{E}$ $=\frac{1}{2}\partial_{i}\mathbf{n}\cdot\partial_{i}\mathbf{n}+\kappa^{2}+\left(m^{2}-\frac{\kappa^{2}}{2}\right)\big{(}1-(n^{3})^{2}\big{)}$
+$\phantom{=\ }+\begin{cases}\kappa\mathbf{n}\cdot\nabla\times\mathbf{n},&\textrm{Dresselhaus SOC},\\ \kappa\left(\mathbf{n}\cdot\nabla n^{3}-n^{3}\nabla\cdot\mathbf{n}\right),\qquad&\textrm{Rashba SOC}.\end{cases}$ (A.5)
For the gauge field corresponding to the Dresselhaus SOC, the Lagrangian is simply given by eq. (A.1) up to a constant ($\kappa^{2}$) with $12m^{2}-\kappa^{2}=$ 1. The other gauge field gives rise to a seemingly different form of the DM term. It is known that the Dresselhaus-SOC-type DM term gives rise to the Bloch-type DW-Skyrmion, whereas the Rashba-SOC-type DM term gives rise to the Néel-type [46]. On the other hand, in the energy functional (A.1), the difference between the two types of DM term is simply a gauge choice. In fact, a gauge transformation
$A_{i}^{a}$ $\to R^{ab}A_{i}^{b}+\mathrm{i}\operatorname{tr}[\tau^{a}U^{-1}\partial_{i}U],$ (A.6)
$n^{a}$ $\to R^{ab}n^{b},\qquad R^{ab}=\frac{1}{2}\operatorname{tr}[\tau^{a}U\tau^{b}U^{-1}],$ (A.7)
with $U\in\operatorname{SU}(2)$ and $R\in\operatorname{SO}(3)$ leaves the energy functional (A.1) invariant. The following gauge transformation
$U=\begin{pmatrix}e^{-\frac{\mathrm{i}\pi}{4}}&0\\ 0&e^{\frac{\mathrm{i}\pi}{4}}\end{pmatrix},\qquad R^{ab}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&1\end{pmatrix},$ (A.8)
changes the gauge field from the Dresselhaus SOC type to the Rashba SOC type and in turn $\mathbf{n}=(n^{1},n^{2},n^{3})\to(-n^{2},n^{1},n^{3})$. This transformation maps a Bloch DW to a Néel DW.